# A Note on Sample Complexity of Interactive Imitation Learning with Log Loss

### v1

@@ Abstract
Imitation learning (IL) is a general paradigm for learning from experts in sequential decision-making problems. Recent advancements in IL [4] have shown that offline imitation learning, specifically Behavior Cloning (BC) with log loss, is minimax optimal. Meanwhile, its interactive counterpart, DAgger, is shown to suffer from suboptimal sample complexity. In this note, we focus on realizable deterministic expert and revisit interactive imitation learning, particularly DAgger with log loss. We demonstrate: 1. A one-sample-per-round DAgger variant that outperforms BC in state-wise annotation. 2. Without recoverability assumption, DAgger with first-step mixture policies matches the performance of BC. Along the analysis, we introduce a new notion of decoupled Hellinger distance that separates state and action sequences, which can be of independent interest.
Yichen Li Chicheng Zhang
University of Arizona University of Arizona
yichenl@arizona.edu chichengz@cs.arizona.edu
@@ 1 Introduction
Imitation learning (IL) is a general paradigm for learning to make sequential decisions. Given offline demonstrations or an interactive expert, the goal is to learn policy that matches the performance of the expert. Two prominent frameworks in IL are offline IL [13] and interactive IL [15]. A popular method of offline IL, behavior cloning (BC), treats IL as a supervised learning problem by collecting trajectories from the expert and learning a policy that maps states to actions. However, it is well known that this may result in compounding error [12], where the learner makes an initial mistake and fails to recover form consecutive unseen states.
To address this challenge, interactive IL [15] introduces training-time expert interaction and utilizes it by iteratively collecting states visited by the learner and querying the expert for action annotations. Interactive IL has been previously known to outperform offline IL under the recoverability assumption [15, 14, 20, 9], where the expert can recover from action deviations with low extra cost.
Recent advancement [4] raises a critical question about the utility of interaction in imitation learning. Specifically, they show that offline IL achieves minimax-optimal sample complexity for both interactive and offline IL, where the sample complexity is in terms of the number of full trajectories annotated. The benefits of interaction for imitation learning highlighted in [4] are restricted to specific scenarios. In this note, we ask the following question:
Can interactive imitation learning be more sample-efficient than offline imitation learning in general?
To answer this, we revisit and conduct refined analysis of (variants) of existing interactive IL algorithms. First, we demonstrate that, when the expert provides state-wise annotations, under a low-recoverability setting, interactive IL can achieve significantly improved sample complexity compared to offline IL. Second, when the expert provides trajectory-wise annotations, we show that without the recoverability assumption, an interactive IL algorithm has sample complexity no worse than that of behavior cloning, contrary to prior claims [4]. Together, these results highlight the potential benefits of interaction in imitation learning and calls for a systematic understanding on imitation learning with state-wise annotations.
@@ 2 Preliminaries
Define $[n]:=\left\{1,\ldots,n\right\}$. Denote by $\Delta(\mathcal{X})$ the set of probability distributions over a finite set $\mathcal{X}$. For $u\in\Delta(\mathcal{X})$ and $x\in\mathcal{X}$, we denote by $u(x)$ the $x$-th coordinate of $u$ and $e_{x}$ the delta mass on $x$. We use the shorthand $x_{1:H}$ to represent the sequence $\{x_{i}\}_{i=1}^{n}$.
Episodic Markov decision process and policy interaction.
@@
Additional policy related notations.

### v2

# Interactive and Hybrid Imitation Learning: Provably Beating Behavior Cloning

@@ Abstract
Imitation learning (IL) is a paradigm for learning sequential decision-making policies from experts, leveraging offline demonstrations, interactive annotations, or both. Recent advances show that when annotation cost is tallied per trajectory, Behavior Cloning (BC)—which relies solely on offline demonstrations—cannot be improved in general, leaving limited conditions for interactive methods such as DAgger to help. We revisit this conclusion and prove that when the annotation cost is measured per state, algorithms using interactive annotations can provably outperform BC. Specifically: (1) we show that Stagger, a one‑sample‑per‑round variant of DAgger, provably beats BC under low-recovery-cost settings; (2) we initiate the study of hybrid IL where the agent learns from offline demonstrations and interactive annotations. We propose Warm-Stagger whose learning guarantee is not much worse than using either data source alone. Furthermore, motivated by compounding error and cold‑start problem in imitation learning practice, we give an MDP example in which Warm-Stagger has significant better annotation cost; (3) experiments on MuJoCo continuous‑control tasks confirm that, with modest cost ratio between interactive and offline annotations, interactive and hybrid approaches consistently outperform BC. To the best of our knowledge, our work is the first to highlight the benefit of state‑wise interactive annotation and hybrid feedback in imitation learning.
@@ 1 Introduction
Imitation learning, or learning from demonstrations, is a widely applied paradigm for learning sequential decision-making policies pomerleau1988alvinn; bojarski2016end; bansal2018chauffeurnet. In many applications, it offers a preferable alternative to reinforcement learning, as it bypasses the need for carefully designed reward functions and avoids costly exploration osa2018algorithmic; sun2017deeply.
Two prominent data collection regimes exist in imitation learning: offline and interactive. In offline imitation learning, expert demonstration data in the format of trajectories is collected ahead of time, which is a non-adaptive process that is easy to maintain. In contrast, in interactive imitation learning, the learner is allowed to query the expert for annotations in an adaptive manner ross2011reduction; ross2014reinforcement; sun2017deeply. The most basic and well-known approach for offline imitation learning is Behavior Cloning ross2010efficient; foster2024behavior, which casts the policy learning problem as a supervised learning problem that learns to predict expert actions from states. Although simple and easy to implement, offline imitation learning has the drawback that the quality of the data can be limited pomerleau1988alvinn. As a result, the trained model can well suffer from compounding error, where imperfect imitation leads the learned policy to enter unseen states, resulting in a compounding sequence of mistakes. In contrast, in interactive imitation learning, the learner maintains a learned policy over time, with the demonstrating experts providing corrective feedback on-policy, which enables targeted collection of demonstrations and improves sample efficiency.
Recent work foster2024behavior, via a sharp theoretical analysis of Behavior Cloning, shows that the sample efficiency of Behavior Cloning cannot be improved in general when measuring using the number of trajectories annotated. Interactive methods like DAgger ross2010efficient can enjoy sample complexity benefits, but so far the benefits are only exhibited in limited examples, with the most general ones in the tabular setting rajaraman2021value. This leaves open the question:
Can interaction provide sample efficiency benefit for imitation learning under a broad range of settings, especially with function approximation?
In this paper, we make progress towards this question, with a focus on the deterministically realizable setting (i.e. the expert policy $\pi^{E}$ is deterministic and is in the learner’s policy class $\mathcal{B}$). Specifically, we make the following contributions:
  1. 1.
    Motivated by the costly nature of interactive labeling on entire trajectories laskey2017comparing; mandlekar2020human, we propose to measure the cost of annotation using the number of states annotated by the demonstrating expert. We propose a general state-wise interactive imitation learning algorithm, Stagger, and show that as long as the expert can recover from mistakes at low cost in the environment ross2011reduction, it significantly improves over Behavior Cloning in terms of its number of state-wise demonstrations required.
  2. 2.
    Motivated by practical imitation learning applications where sets of offline demonstration data are readily available, we study hybrid imitation learning, where the learning agent has the additional ability to query the demonstration expert interactively to improve its performance. We design a hybrid imitation learning algorithm, Warm-Stagger, and prove that its policy optimality guarantee is not much worse than using either of the data sources alone.
  3. 3.
    Inspired by compounding error pomerleau1988alvinn and cold start problem lin2020autoregressive; oetomo2023cutting, two practical challenges in imitation learning, we provide an MDP example, for which we show hybrid imitation learning can achieve strict sample complexity savings over using either source alone, and provide simulation results that verify this theoretical claim.
  4. 4.
    We conduct experiments in MuJoCo continuous control tasks and show that if the cost of state-wise interactive demonstration is not much higher than its offline counterpart, interactive algorithms can enjoy a better cost efficiency than Behavior Cloning. Under some cost regimes and some environments, hybrid imitation learning can outperform approaches that use either source alone.
@@ 2 Preliminaries
Define $n]:=\left\{1,\ldots,n\right\}$. Denote by $\Delta(\mathcal{X})$ the set of probability distributions over a set $\mathcal{X}$. For $u\in\Delta(\mathcal{X})$ and $x\in\mathcal{X}$, we denote by $u(x)$ the $x$-th coordinate of $u$ and $e_{x}$ the delta mass on $x$. We use the shorthand $x_{1:n}$ to represent the sequence $\{x_{i}\}_{i=1}^{n}$. We will frequently use the Hellinger distance to measure the difference between two distributions: $D_{H}^{2}(\mathbb{P},\mathbb{Q})=\int(\sqrt{\tfrac{d\mathbb{P}}{d\omega}}-\sqrt{\tfrac{d\mathbb{Q}}{d\omega}})^{2}d\omega$, where $\mathbb{P}$ and $\mathbb{Q}$ share a dominating measure $\omega$.
Episodic Markov decision process and agent-environment interaction.
An episodic MDP $\mathcal{M}$ is defined as a tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R},H)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:=\{P_{h}:{\mathcal{S}}\times\mathcal{A}\rightarrow\Delta({\mathcal{S}})\}_{h=1}^{H}$ denotes the transition dynamics, $\mathcal{R}:{\mathcal{S}}\times\mathcal{A}\rightarrow\Delta([0,1])$ denotes the reward distribution, and $H$ denotes episode length. Given a stationary policy $\pi:{\mathcal{S}}\rightarrow\Delta(\mathcal{A})$, we use $\pi(\cdot|s)$ to denote the action distribution of $\pi$ on $s$. A policy induces a distribution over trajectories $\tau=(s_{h},a_{h},r_{h})_{h=1}^{H}$ by first drawing the initial state $s_{1}\sim P_{0}(\varnothing)$, and then iteratively taking actions $a_{h}\sim\pi(\cdot\mid s_{h})$, receiving rewards $r_{h}\sim\mathcal{R}(s_{h},a_{h})$, and transitioning to the next state $s_{h+1}\sim P_{h}(s_{h},a_{h})$ (except at step $H$, where $P_{H}=\varnothing$). Let $\mathbb{E}^{\pi}$ and $\mathbb{P}^{\pi}$ denote expectation and probability law for $(s_{h},a_{h},r_{h})_{h=1}^{H}$ induced by $\pi$ and $\mathcal{M}$. Given $\pi$, denote by $d^{\pi}(s):=\frac{1}{H}\mathbb{P}^{\pi}(s_{h}=s)$ its state visitation distribution. The expected return of policy $\pi$ is defined as $J(\pi):=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}r_{h}\right],$ and the value functions of $\pi$ are given by $V_{h}^{\pi}(s):=\mathbb{E}^{\pi}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}\,|\,s_{h}=s\right]$, and $Q_{h}^{\pi}(s,a):=\mathbb{E}^{\pi}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}\,|\,s_{h}=s,a_{h}=a\right].$ If for policy $\pi$, step $h$, and state $s$, $\pi(\cdot\mid s)$ is the delta-mass on an action, we also sometimes slightly abuse the notation and let $\pi(s)$ denote that action.
Additional policy-related notations.
Throughout, we assume the access to an Markovian policy class $\mathcal{B}$ of finite size $B$, which contains the deterministic expert demonstrator policy $\pi^{\mathrm{E}}:{\mathcal{S}}\rightarrow\Delta(\mathcal{A})$. A $(\text{MDP},\text{Expert})$ pair $(\mathcal{M},\pi^{\mathrm{E}})$ is said to be $\mu$-recoverable if for all $h\in[H]$, $s\in{\mathcal{S}}$ and $a\in\mathcal{A}$, $Q_{h}^{\pi^{\mathrm{E}}}(s,a)-V_{h}^{\pi^{\mathrm{E}}}(s)\leq\mu$. Additionally, we assume normalized return foster2024behavior, where for any trajectory $(s_{h},a_{h},r_{h})_{h=1}^{H}$, $\sum_{h=1}^{H}r_{h}\in[0,R]$. Throughout this paper, we make the assumption that our imitation learning problem is deterministically realizable:
Assumption 1 (Deterministic Realizability).
The expert policy $\pi^{\mathrm{E}}$ is deterministic and is contained in the learner’s policy class $\mathcal{B}$.
In our algorithm and analysis, we frequently use the following “convexification” of policy class $\mathcal{B}$:
Definition 1 (Each-step Mixing of $\mathcal{B}$).
$\bar{\Pi}_{\mathcal{B}}=\{\bar{\pi}_{u,h}(a|s):=\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a|s):u\in\Delta(\mathcal{B})\}.$
An each-step mixing policy $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$ can be executed by drawing $\pi\sim u$ freshly-at-random at each step $h\in[H]$ and takes action $a\sim\pi(\cdot|s)$ (e.g. li2022efficient; li2023agnostic). Observe that $\bar{\pi}_{u}$ is a Markovian policy.
Offline imitation learning and Behavior Cloning.
In offline imitation learning, the agent is given a collection of expert trajectories $\mathcal{D}=\left\{\tau_{1},\ldots,\tau_{N_{\mathrm{off}}}\right\}$, where $\tau_{i}=(s_{i,h},a_{i,h})_{h=1}^{H}$ is the $i$-th (reward-free) trajectory, all of which are drawn iid from the trajectory distribution of expert policy $\pi^{\mathrm{E}}$. Behavior Cloning finds a policy $\pi\in\mathcal{B}$ that minimizes its log loss on expert’s actions on the seen states, i.e.,
$\hat{\pi}=\mathop{\mathrm{argmin}}_{\pi\in\mathcal{B}}\sum_{i=1}^{N_{\mathrm{off}}}\sum_{h=1}^{H}\log\frac{1}{\pi(a_{i,h}\mid s_{i,h})}.$
Recent result of foster2024behavior establishes a horizon-independent analysis of Behavior Cloning, which we recall its guarantees here:
Theorem 2 (Guarantee of BC foster2024behavior).
Suppose Assumption 1 holds, then with probability $1-\delta$, the policy returned by BC $\hat{\pi}$ satisfies:
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq\tilde{O}\left(\frac{R\log B}{N_{\mathrm{off}}}\right).$
Interactive imitation learning protocol.
In interactive IL, the learner has the ability to query the demonstration expert interactively. A first way to model interaction with expert is through a trajectory-wise demonstration oracle $\mathcal{O}^{\mathrm{Traj}}$ ross2011reduction; foster2024behavior: given a state sequence $(s_{h})_{h=1}^{H}$, return $(a_{h})_{h=1}^{H}$ such that $a_{h}=\pi^{\mathrm{E}}(s_{h})$ for all $h$. Subsequent works have considered modeling the expert as a state-wise demonstration oracle judah2014active; brantley2019disagreement; nguyen2020active; sekhari2024selective $\mathcal{O}^{\mathrm{State}}$: given a state $s_{h}$ and step $h$, return $a_{h}=\pi^{\mathrm{E}}(s_{h})$. We consider the learner interacting with the environment and demonstration oracles using the following protocol:
  • For $i=1,2,\dots$
    • Select policy $\pi^{i}$ and roll it out in $\mathcal{M}$, observing a reward-free trajectory $(s_{1},a_{1},\ldots,s_{H},a_{H})$.
    • Query the available oracle(s) to obtain expert annotations.
@@
In practice, we expect the cost of querying $\mathcal{O}^{\mathrm{Traj}}$ to be higher than that of collecting a single offline expert trajectory laskey2017comparing. Since $H$ queries to $\mathcal{O}^{\mathrm{State}}$ can simulate one query to $\mathcal{O}^{\mathrm{Traj}}$, the cost of a single $\mathcal{O}^{\mathrm{State}}$ query should be at least $\frac{1}{H}$ the cost of $\mathcal{O}^{\mathrm{Traj}}$. Consequently, we also expect one $\mathcal{O}^{\mathrm{State}}$ query to be more expensive than obtaining an additional offline (state, expert action) pair. We denote the ratio between these two costs as $C$, where $C\geq 1$ is an application-dependent constant.  [^1] For practical settings such as human-in-the-loop learning with expert interventions mandlekar2020human; spencer2020learning, obtaining a short segment of corrective demonstrations may be cheaper than querying $\mathcal{O}^{\mathrm{State}}$ for each state therein. Here, we focus on a simplified setting and leave detailed cost modeling for such settings as interesting future work.
3 State-wise Annotation in Interactive Imitation Learning
Recent work foster2024behavior on refined analysis of Behavior Cloning (BC) casts doubt on the utility of interaction in imitation learning: when measuring sample complexity in the number of trajectories annotated, BC is minimax optimal even among interactive algorithms (foster2024behavior, Corollary 2.1 and Theorem 2.2). Although benefits of interactive approaches have been shown in specific examples, progresses so far have been sparse foster2024behavior; rajaraman2021value, with the most general results in the less-practical tabular setting rajaraman2021value. In this section, we show that interaction benefits imitation learning in a fairly general sense: when measuring sample complexity using the number of state-wise annotations, we design an interactive algorithm with sample complexity better than BC, as long as the expert policy has a low recovering cost $\mu$ in the environment.
Algorithm 1 Stagger: DAgger with State-wise annotation oracle
1:Input: MDP $\mathcal{M}$, state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$ with query budget $N_{\mathrm{int}}$, Markovian policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$.
2:for $n=1,\dots,N_{\mathrm{int}}$ do
3:  Query $\mathbb{A}$ and receive $\pi^{n}$.
4:  Execute $\pi^{n}$ and sample $s^{n}\sim d^{\pi^{n}}$. Query $\mathcal{O}^{\mathrm{State}}$ for $a^{*,n}=\pi^{\mathrm{E}}(s^{n})$.
5:  Update $\mathbb{A}$ with loss function
$\ell^{n}(\pi):=\log\left(\frac{1}{\pi(a^{*,n}|s^{n})}\right).$ (1)
6:end for
7: Output $\hat{\pi}$, a first-step uniform mixture of $\left\{\pi^{n}\right\}_{n=1}^{N_{\mathrm{int}}}$.
3.1 Interactive IL Enables Improved Sample Complexity with State-wise Annotations
Our algorithm Stagger (short for State-wise DAgger), namely Algorithm 1, interacts with the demonstration expert using a state-wise annotation oracle $\mathcal{O}^{\mathrm{State}}$. Similar to the original DAgger ross2011reduction, it requires base policy class $\mathcal{B}$ and reduces interactive imitation learning to no-regret online learning. At round $n$, it rolls out the current policy $\pi^{n}$ obtained from an online learning oracle $\mathbb{A}$ and samples $s^{n}$ from $d^{\pi^{n}}$. A classical example of $\mathbb{A}$ is the exponential weight algorithm that chooses policies from $\bar{\Pi}_{\mathcal{B}}$ (cesa2006prediction; see Proposition 38 in Appendix F). It then queries $\mathcal{O}^{\mathrm{State}}$ to get expert action $a^{*,n}$ and updates $\mathbb{A}$ with loss function $\ell^{n}(\pi)$ induced by this new example (Eq. (1)). The final policy $\hat{\pi}$ is returned as a uniform first-step mixture of the historical policies $\{\pi^{n}\}^{N_{\mathrm{int}}}_{n=1}$, i.e., sample one $\pi^{n}$ uniformly at random and execute it for the episode. In contrast to the DAgger variant analyzed in foster2024behavior, which trains a distinct policy at each step—yielding $H$ policies in total—and employs trajectory-level annotations, our algorithm utilizes parameter sharing and uses state-wise annotations.
We show the following performance guarantee of Algorithm 1 with $\mathbb{A}$ instantiated as the exponential weight algorithm:
Theorem 3.
Suppose Stagger is run with a state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$, an MDP $\mathcal{M}$ where $(\mathcal{M},\pi^{\mathrm{E}})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that deterministic realizability (Assumption 1) holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm with decision space $\Delta(\mathcal{B})$ and returns each-step mixing policies $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$. Then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq\mu H\cdot\frac{\log(B)+2\log(1/\delta)}{N_{\mathrm{int}}}.$
Theorem 3 shows that Stagger returns a policy of suboptimality $O(\frac{\mu H\log B}{N_{\mathrm{int}}})$ using $N_{\mathrm{int}}$ interactive state-wise annotations from the expert. In comparison, with the cost of $N_{\mathrm{int}}$ state-wise annotations, one can obtain $\frac{CN_{\mathrm{int}}}{H}$ trajectory-wise annotations; foster2024behavior’s analysis shows that Behavior Cloning with this number of trajectories from $\pi^{\mathrm{E}}$ returns a policy of suboptimality $O(\frac{RH\log B}{CN_{\mathrm{int}}})$; recall Theorem 2. Thus, if $C\ll\frac{R}{\mu}$, Algorithm 1 has a better cost-efficiency guarantee than Behavior Cloning.
We now sketch the proof of Theorem 3. In line with foster2024behavior, we define the online, on-policy state-wise estimation error as
$\operatorname{OnEst}^{\mathrm{State}}_{N}:=\sum_{n=1}^{N}\mathbb{E}_{s\sim d^{\pi^{n}}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s),\pi^{\mathrm{E}}(\cdot\mid s))\right].$
The proof proceeds by bounding this error and translating it to the performance difference between $\hat{\pi}$ and $\pi^{\mathrm{E}}$. While our definition of estimation error is similar to (foster2024behavior, Appendix C.2), their definition requires all $H$ states per trajectory, while ours depends on the distribution over a state sampled uniformly from the rollout of policy $\pi^{n}$. This enables each labeled state to serve as immediate online feedback, fully utilizing the adaptivity of online learning. In contrast, trajectory-wise annotations may cause the online learning oracle to operate under delayed feedback joulani2013online; wan2022online, which incurs a fundamental extra factor of $H$ in state-wise sample complexity compared to our approach.
3.2 Experimental Comparison
We conduct a simple simulation study comparing the sample efficiency of log-loss Behavior Cloning foster2024behavior and Stagger in four MuJoCo todorov2012mujoco; brockman2016openai continuous control tasks with $H=1000$ and pretrained deterministic MLP experts schulman2015trust; schulman2017proximal. Considering MuJoCo’s low sensitivity to horizon length foster2024behavior, we reveal expert states one by one along consecutive trajectories for BC to allow fine-grained state-wise sample complexity comparison, while Stagger queries exactly one state per iteration by sampling from the latest policy’s rollout and updating immediately with the expert’s annotation. In Stagger, we implement the online learning oracle $\mathbb{A}$ so that it outputs a policy that approximately minimizes the log loss. In addition to log loss, we also include results with online learning oracle minimizing historical examples’ total square loss in Appendix G.2. We defer other implementation details to Appendix G.
Figure 1 shows the performance of the learned policy as a function of the number of state-wise annotations. When each interactive state-wise annotation has the same cost as an offline (state, expert action) pair ($C=1$), Stagger has superior and more stable performance than Behavior Cloning. For a given target performance (e.g., near expert-level), Stagger often requires significantly fewer state-wise annotations than BC—especially on harder tasks—though the gains are less pronounced on easier ones like Ant and Hopper. To highlight sample efficiency, we plot Stagger using only half the annotation budget of BC; despite this, it still matches or surpasses BC on several tasks, suggesting meaningful benefits from interaction when $C$ is small (e.g., $C=3$ for Walker).
Figure 1: State-wise sample complexity comparison between Behavior Cloning and Stagger. Shaded areas show the 10th–90th percentile bootstrap confidence intervals diciccio1996bootstrap over 10 runs. Stagger matches or exceeds BC with $50\%$ fewer annotations, achieving better state-wise annotation efficiency.
4 Hybrid Imitation Learning: Combining Offline Trajectory-wise and Interactive State-wise Annotations
Practical deployments of imitation learning systems often learn simultaneously from offline and interactive feedback modalities kelly2019hg; hoque2021thriftydagger: for example, in autonomous driving zhang2017query; badue2021self; zhao2023real, the learner has access to some offline expert demonstrations to start with, and also receives interactive expert demonstration feedback in trajectory segments for subsequent finetuning. Motivated by this practice, we formulate the following problem:
Hybrid Imitation Learning (HyIL): Problem Setup.
The learner has access to two complementary sources of expert supervision:
  • $N_{\mathrm{off}}$ offline expert trajectories $D_{\mathrm{off}}=\left\{(s_{i,h},a_{i,h})_{h=1}^{H},i\in[N_{\mathrm{off}}]\right\}$, sampled i.i.d. from rolling out $\pi^{E}$ in $\mathcal{M}$;
  • A state-wise annotation oracle $\mathcal{O}^{\mathrm{State}}$ that can be queried interactively up to $N_{\mathrm{int}}$ times.